What is the perimeter of a square whose diagonal is 5 times the square root of 2?
Answer by nyc_function(2741)

Use the Pythagorean Theorem: a^ + b^2 = c^2
We do not know the measure of one side of this square.
The sides of any square is the same all around.
In the theorem above, a^2 and b^2 represent two sides of your square.
However, c^2 represents 5(sqrt{2}).
NOTE: 5(sqrt{2}) is read: "five times the square root of two."
Let x = one side of your square.
Then we have x^2 + x^2 = [5(sqrt{2})]^2
2x^2 = 50
x^2 = 50/2
x^2 = 25
x = 5
The perimeter of any square can found using the formula P = 4 times side.
I just found one side of your square to be 5.
Then P = 4 times 5 or 20.
The perimeter is 20.
NOTE: You can also add all four sides to find the perimeter of your
square like this: 5 + 5 + 5 + 5 = 20.
nycfunction@yahoo.com
Guido